# Adhesion Energy of Phosphorene on different pristine and oxidized metallic substrates

### v1

@@ Abstract
Black phosphorus and its single-layer constituent, phosphorene, have emerged as promising two-dimensional materials with remarkable tribological properties. However, recent experimental investigations reveal complex substrate-dependent behavior that affects their lubricating capabilities. This computational study employs density functional theory calculations to quantify the adhesion energy of both pristine and oxidized phosphorene monolayers on various metallic substrates (aluminum, copper, iron, and chromium) and their corresponding oxides ($\mathrm{Al_{2}O_{3}}$, $\mathrm{Cu_{2}O}$, $\mathrm{Fe_{2}O_{3}}$, and $\mathrm{Cr_{2}O_{3}}$), correlating these fundamental interactions with experimentally observed tribological performance. Results demonstrate that oxidized phosphorene exhibits greater stability than its pristine counterpart and shows higher adhesion to all substrates, attributed to favorable interactions between oxygen non-bonding states and substrate empty states. Adhesion is systematically more favorable on pristine metals than on their corresponding oxides, with chromium and iron showing particularly strong interactions due to partially filled 3d orbitals. This result is consistent with the coefficient of friction decrease observed in tribological experiments after scratching the substrate, thus removing the outermost oxide layer. Charge redistribution and electronic structure analyses reveal system-dependent interfacial bonding characteristics, with some configurations inducing metallic character in phosphorene. These findings provide fundamental insights into substrate-dependent black phosphorus lubricating properties, highlighting the key role of layer-substrate adhesion.
@@ 1 Introduction
Phosphorene, defined as a single layer exfoliated from bulk BP, represents a particularly important member of this material class. Its unique electronic properties have attracted considerable attention, particularly its direct bandgap of 0.3 eV [kou2015phosphorene, carvalho2016phosphorene], which can be systematically tuned by varying the number of layers [low2014tunable] or applying in-plane strain [rodin2014strain]. Beyond its electronic characteristics, phosphorene exhibits remarkable mechanical and tribological properties attributed to its anisotropic Young’s modulus [tao2015mechanical] and its exceptional ability to achieve ultralow friction when employed as a lubricant additive in aqueous solutions [wang2018superlubricity].
The superior tribological performance of phosphorene as both filler and composite material [wang2018superlubricity, cui2017atomic, lv2018self, wang2018black] originates from the ability to reduce the surface energy of the substrate [restuccia2016tribochemistry, marchetto2017surface] and the weak van der Waals interactions between layers [philippon2007experimental, de2015tribochemistry], which facilitate interlayer sliding and reduce friction. However, recent experimental investigations have revealed that BP’s solid lubrication performance exhibits complex substrate-dependent behavior that challenges conventional assumptions about its lubricating capabilities [vezzelli2024different].
From a structural perspective, BP possesses a distinctive atomic arrangement that governs its tribological properties. Each phosphorus atom contains five valence electrons in a $3s^{2}3p^{3}$ configuration, forming four $sp^{3}$ hybridized orbitals. Three orbitals participate in bonding with adjacent phosphorus atoms, while the fourth hosts a lone electron pair, resulting in a tetrahedral geometry that produces a characteristic puckered honeycomb lattice structure. This arrangement exhibits pronounced structural anisotropy: the material displays zigzag and armchair configurations along the x- and y-directions, respectively, while exhibiting a hexagonal structure along the z-direction. Each phosphorene layer comprises two atomic sub-layers connected by two distinct types of P-P bonds: intra-planar bonds within each sub-layer and weaker inter-planar bonds between sub-layers, as illustrated in the Supplementary Material [losi2020superlubricity, benini2023interaction, zhang2021semiconducting, boddula2020black, losi2023modeling].
@@
The substrate-dependent tribological behavior of BP was systematically investigated in the study ”A Different Perspective on the Solid Lubrication Performance of Black Phosphorus: Friend or Foe?” [vezzelli2024different], which demonstrated that BP’s performance varies significantly with substrate material and contact conditions. That work identified adhesion between BP and the substrate as a critical factor determining its effectiveness as a solid lubricant. Building upon these experimental findings, the present computational study extends this analysis by quantifying the adhesion energy of BP on identical metal substrates (steel, aluminum, copper, and iron) through density functional theory (DFT) calculations.
Figure 1: Comparison between pristine and oxidized phosphorene structures. (a) Bottom view and (c) side view of pristine phosphorene. (b) Bottom view and (d) side view of oxidized phosphorene with 25% of oxygen coverage.
@@ 2 Computational details
Since DFT faces significant difficulties in accurately modeling the electronic structure and magnetic properties of transition-metal oxides, the DFT+U method was employed to better capture the strong localization of the d orbitals in these systems, accounting for the on-site Coulomb interactions [naveas2023first]. The Hubbard parameters were determined empirically, with iterative adjustments made to optimize structural parameters and band gap values, ensuring an accurate description of the transition-metal oxides. The values used were 4.8 eV for the 3d orbitals of $\mathrm{Fe_{2}O_{3}}$, 7.0 eV for the 3d orbitals of $\mathrm{Cr_{2}O_{3}}$, and 11.0 eV for the 3d and 2p orbitals of $\mathrm{Cu_{2}O}$. The Hubbard plots for the bulk materials are shown in the Supplementary Material.
Table 1 presents a comprehensive overview of the computational details used for bulk and substrate calculations. It includes the optimized bulk lattice parameters, calculated surface energies corresponding to the most stable surfaces and terminations, and the number of substrate layers used.
The optimized structural parameters and surface energies obtained in this study were validated against literature data. For pristine phosphorene, the calculated lattice parameters and surface energy are consistent with ref. [dhanabalan2017emerging]. The surface energies of Al(111), Cu(111), and Fe(110) agree with previous DFT calculations [jain2013commentary, kiejna2001first, vitos1998surface]. For $\mathrm{Al_{2}O_{3}}$(0001), both Al-terminated and O-terminated configurations were benchmarked against refs. [chiang2016dft, ramogayana2021density, sun2006structure]. The $\mathrm{Cu_{2}O}$(111) O-terminated surface structure and energetics are consistent with refs. [zhang2018atomistic, kokalj2015density, gao2020surface], while $\mathrm{Fe_{2}O_{3}}$(0001) Fe-terminated and O-terminated surfaces were compared with refs. [huang2016surface, nguyen2013water, rohrbach2004ab]. The Cr(110) surface energy matches literature values [jain2013commentary, vitos1998surface], and $\mathrm{Cr_{2}O_{3}}$(0001) Cr-terminated surface with refs. [sun2006structure, rohrbach2004ab, costa2009ab]. The number of layers for each substrate ensures surface energy convergence as reported in the corresponding references.
@@
For this study, calculations focused solely on monolayer phosphorene, as the primary interest was in determining adhesion energy. This approach optimized computational efficiency and resource utilization. Notably, other computational studies have employed bilayer phosphorus models only when investigating metal interfaces [losi2023modeling] or conducting molecular dynamics (MD) simulations [benini2023interaction].
For oxidized phosphorene, the bulk structure was constructed by replicating the most stable oxidized monolayer configuration along the out-of-plane direction. The bulk total energy ($E_{\mathrm{bulk}}$) used in Equation 1 for surface energy calculations was obtained from this multilayer oxidized structure, ensuring consistency with the oxidation pattern of the monolayer system.
@@
For the metal oxide substrates, different surface terminations were investigated by cleaving the structures along the z-axis to expose different outermost atomic layers. The notation used throughout this work indicates the chemical nature of the topmost layer: ”Al-term” refers to an aluminum-terminated surface, ”O-term” to an oxygen-terminated surface, ”Fe-term” to an iron-terminated surface, ”Cr-term” to a chromium-terminated surface, and ”Cu-term” to a copper-terminated surface. The relaxed geometries of these different surface terminations are shown in the Supplementary Material.
@@
where $E_{\mathrm{sub}}$ and $E_{\mathrm{bulk}}$ are the total energies of the substrate and of the bulk, respectively, n is the ratio between the number of atoms in the substrate and in the bulk, and A is the area of the simulation cell. For phosphorene, the monolayer was considered for the calculation of $E_{\mathrm{sub}}$.
@@
To quantify the structural differences between the interacting structures in the supercell and their isolated counterparts, the Root Mean Square Deviation (RMSD) was employed through a custom Python script. RMSD is a widely adopted metric for expressing structural similarity by calculating the average deviation between equivalent atoms in two structures. While commonly used in protein structure analysis and MD simulations, it is also applicable in DFT studies [debefve2021systematic, sharma2014accuracy]. The RMSD is defined as:
$\mathrm{RMSD}=\sqrt{\frac{\sum_{i}d_{i}^{2}}{n}}$ (3)
@@
where $d_{i}$ represents the distance between each of the n pairs of equivalent atoms in two optimally superimposed structures. The structures were aligned to minimize the effects of translation and rotation. An RMSD value of 0 Å indicates identical structures, with increasing values reflecting greater structural differences [carugo2001normalized]. RMSD values are considered reliable indicators of variability when applied to highly similar structures, providing a quantitative measure of geometrical resemblance [carugo2001normalized]. In this study, visual inspection of the phosphorene-substrate interactions revealed noticeable but not excessive distortions. Therefore, RMSD evaluation was considered an appropriate approximation for quantifying the degree of structural distortion.
To analyze the behavior of electronic charge resulting from the interaction between phosphorene and the substrate, the total electron charge redistribution per surface unit ($Q_{red}$) was computed by first subtracting the charge densities of the isolated components from the charge density of the combined system to obtain the charge density difference, $\Delta\rho(x,y,z)$. The resulting $\Delta\rho(x,y,z)$ was then integrated along the xy-plane to obtain the planar average charge density redistribution, $\Delta\rho(z)$. Finally, $Q_{red}$ was obtained by integrating the absolute value of $\Delta\rho(z)$ along the z-axis and divided by the supercell area. This procedure was adapted from ref. [wolloch2018interfacial].
@@ 3 Results and discussion
For oxide substrates, the adhesion energies are generally less favorable (less negative values) compared to their corresponding pristine metals. This can be attributed to the reduced availability of electronic charge for bonding, as oxides typically have fewer delocalized electrons than metals. Among the oxides, $\mathrm{Fe_{2}O_{3}}$ shows the strongest adhesion, particularly with oxidized phosphorene and oxygen termination. $\mathrm{Cu_{2}O}$ exhibits better adhesion with pristine phosphorene, while $\mathrm{Cr_{2}O_{3}}$ shows similar values for both forms. Notably, while $\mathrm{Al_{2}O_{3}}$ demonstrates comparable adhesion to other oxides with pristine phosphorene, its interaction with oxidized phosphorene is extremely weak.
The relationship between the calculated adhesion energies and the experimental tribological performance reported in Ref. [vezzelli2024different] reveals important insights.
In the experiments, the best results in terms of coefficient of friction (CoF) reduction were observed for scratched iron, where the superficial oxide layer was most likely removed, while non-scratched iron exhibited poor performance. This suggests that phosphorene-substrate adhesion plays a crucial role in enhancing the lubricating properties of BP, being significantly higher on pristine iron than on oxidized iron (Figure 3). This observation is consistent with findings for other layered materials used as solid lubricants, such as MXenes [marquis2025mxene], and can be explained by considering that higher adhesion of the layers to the substrate prevents their peeling off during the rubbing process. Indeed, BP was found to reduce the CoF of non-scratched aluminum substrate only at the beginning of the tribological test, while after rubbing it reached the value measured for the unlubricated sample [vezzelli2024different], in agreement with the low phosphorene-$\mathrm{Al_{2}O_{3}}$ adhesion calculated in this work (Figure 3). The low phosphorene-$\mathrm{Cu_{2}O}$ adhesion may suggest an irrelevant effect of BP in reducing the CoF of copper, as observed for aluminum. However, the experimental results indicate an even worse effect of BP on copper, which was interpreted as third-body wear effect due to the poor adhesion.
Figure 4 presents the evaluation of RMSD for phosphorene-substrate interactions. Analysis of the data reveals a weak correlation between RMSD values and adhesion energies. In general, oxidized phosphorene exhibits higher RMSD values compared to its pristine counterpart, aligning with its more favorable adhesion energies. This suggests that oxidized phosphorene undergoes greater structural distortion due to the more favorable interaction. However, the pattern of increased distortion observed for oxidized phosphorene does not extend to the corresponding substrate surfaces. This observation implies that the structural changes are predominantly confined to the phosphorene layer, with minimal deformation occurring in the substrate materials. Notably, $\mathrm{Cu_{2}O}$ stands as an exception to this trend.
@@
Figure 5 presents the overall assessment of charge redistribution ($Q_{red}$) when phosphorene is deposited on the substrate. In agreement with previous studies [losi2023modeling, wolloch2018interfacial], higher interfacial adhesion (i.e., more negative adhesion energies) is associated with higher $Q_{red}$ values. Moreover, $Q_{red}$ follows the same trend as adhesion, increasing from Al to Cu, Fe, and Cr in the case of pristine metals, while it remains very low for oxides, except for O-terminated $\mathrm{Fe_{2}O_{3}}$, where the adhesion is higher than for the other oxides.
Figure 5: Total electron charge redistribution per surface unit ($Q_{red}$) for every phosphorene-substrate system.
Figure 6: $\Delta\rho(z)$ and PDOS analysis for every phosphorene-substrate system.
Figure 7: $\Delta\rho(z)$ and PDOS analysis for every phosphorene-substrate system.
Figures 6 and 7 illustrate the charge analysis, specifically $\Delta\rho(z)$ and PDOS. As shown in these figures, $Q_{red}$ is higher for pristine metals than for their oxidized counterparts. This may be attributed to the fact that pristine metals, with less localized charge compared to metal oxides, have more surface dangling bonds and can more easily form new bonds with phosphorene, increasing the charge redistribution.
Regarding $\Delta\rho(z)$, which does not necessarily correlate with $Q_{red}$ due to its dependence on the supercell area, a greater charge redistribution is generally observed in pristine metals compared to their respective oxides. Overall, charge accumulation is noted in the interfacial region, attributed to the formation of new bonds between phosphorene and the substrate surface.
From the PDOS analysis, an increase in phosphorene p-states within the band gap is observed in some cases after interaction with the substrate (e.g., Figures 6(j) and 6(l)). In these cases, phosphorene exhibits metallic character [pan2016monolayer]. In some instances, metallic behavior emerges only for oxidized phosphorene (e.g., Figures 6(b) and 6(d)), whereas in other cases it is observed for both pristine and oxidized phosphorene (e.g., Figures 6(j) and 6(l)). Conversely, in some systems no p-states appear within the band gap (e.g., Figures 6(b) and 7(b)), thus preserving the insulating character. Moreover, no strong correlation is found between the emergence of metallic behavior in phosphorene and the $Q_{red}$ values, as evidenced by the comparison between Figures 6(j) and 6(l) and the $Q_{red}$ value reported for Cr in Figure 5.

### v2

@@ Abstract
Black phosphorus and its single-layer constituent, phosphorene, have emerged as promising two-dimensional materials with remarkable tribological properties. However, recent experimental investigations revealed that the their lubricating capabilities can change with the substrate. The present computational study employs density functional theory calculations to quantify the adhesion energy of both pristine and oxidized phosphorene monolayers on various metallic substrates (aluminum, copper, iron, and chromium) and their corresponding oxides ($\mathrm{Al_{2}O_{3}}$, $\mathrm{Cu_{2}O}$, $\mathrm{Fe_{2}O_{3}}$, and $\mathrm{Cr_{2}O_{3}}$), correlating these interfacial property with experimentally observed tribological performance. Results demonstrate that oxidized phosphorene presents higher adhesion to all substrates with respect to pristine phosphorene, attributed to favorable interactions between oxygen non-bonding states and substrate empty states. Adhesion is systematically more favorable on pristine metals than on their corresponding oxides, with chromium and iron showing particularly strong interactions due to partially filled 3d orbitals. This result is consistent with the coefficient of friction decrease observed in tribological experiments after scratching the iron substrate, thus removing the outermost oxide layer. Charge redistribution correlates with the adhesion and electronic structure analyses reveal system-dependent interfacial bonding characteristics, with some configurations inducing metallic character in phosphorene. These findings provide fundamental insights into substrate-dependent lubricating properties of black phosphorus, highlighting the key role of layer-substrate adhesion.
@@ 1 Introduction
Phosphorene, a single layer exfoliated from bulk BP, presents unique electronic properties that have attracted considerable attention, owing to its direct bandgap of 0.3 eV [kou2015phosphorene, carvalho2016phosphorene], which can be systematically tuned by varying the number of layers [low2014tunable] or applying in-plane strain [rodin2014strain]. Beyond its electronic characteristics, phosphorene exhibits remarkable mechanical and tribological properties attributed to its anisotropic Young’s modulus [tao2015mechanical] and its exceptional ability to achieve ultralow friction when employed as a lubricant additive in aqueous solutions [wang2018superlubricity].
The superior tribological performance of phosphorene as both filler and composite material [wang2018superlubricity, cui2017atomic, lv2018self, wang2018black] originates from the ability, common to other layered materials, to reduce the surface energy of the substrate [losi2023modeling, restuccia2016tribochemistry, marchetto2017surface] and in the weak interlayer interactions [losi2020superlubricity, philippon2007experimental, de2015tribochemistry], which reduce interfacial adhesion and friction. However, recent experimental investigations have shown that BP performance as solid lubricant is deeply influenced by the substrate [vezzelli2024different].
From a structural perspective, BP possesses a distinctive atomic arrangement that governs its tribological properties. Each phosphorus atom contains five valence electrons in a $3s^{2}3p^{3}$ configuration, forming four $sp^{3}$ hybridized orbitals. Three orbitals participate in bonding with adjacent phosphorus atoms, while the fourth hosts a lone electron pair, resulting in a tetrahedral geometry that produces a characteristic puckered honeycomb lattice structure. This arrangement exhibits pronounced structural anisotropy: the material displays zigzag and armchair configurations along the x- and y-directions, respectively, while exhibiting a hexagonal structure along the z-direction. Each phosphorene layer comprises two atomic sub-layers connected by two distinct types of P-P bonds: intra-planar bonds within each sub-layer and weaker inter-planar bonds between sub-layers, as illustrated in the Supplementary Material (SM) [losi2020superlubricity, benini2023interaction, zhang2021semiconducting, boddula2020black, losi2023modeling].
@@
The substrate-dependent tribological behavior of BP was systematically investigated in the study ”A Different Perspective on the Solid Lubrication Performance of Black Phosphorus: Friend or Foe?” [vezzelli2024different], which demonstrated that BP’s performance varies significantly with substrate material and contact conditions. That work identified adhesion between BP and the substrate as a critical factor determining its effectiveness as a solid lubricant. Building upon these experimental findings, the present computational study extends this analysis by quantifying the adhesion energy of BP on the same metal substrates (steel, aluminum, copper, and iron) through density functional theory (DFT) calculations.
Figure 1: Comparison between pristine and oxidized phosphorene structures. (a) Bottom view and (c) side view of pristine phosphorene. (b) Bottom view and (d) side view of oxidized phosphorene with 25% of oxygen coverage, determined as the most stable configuration for oxidized phosphorene [benini2023interaction].
@@ 2 Computational details
Since DFT faces significant difficulties in accurately modeling the electronic structure and magnetic properties of transition-metal oxides, the DFT+U method was employed to better capture the strong localization of the d orbitals in these systems, accounting for the on-site Coulomb interactions [naveas2023first]. The Hubbard parameters were determined empirically, with iterative adjustments made to optimize structural parameters and band gap values, ensuring an accurate description of the transition-metal oxides. The values used were 4.8 eV for the 3d orbitals of $\mathrm{Fe_{2}O_{3}}$, 7.0 eV for the 3d orbitals of $\mathrm{Cr_{2}O_{3}}$, and 11.0 eV for the 3d and 2p orbitals of $\mathrm{Cu_{2}O}$. The Hubbard plots for the bulk materials are shown as SM.
Table 1 presents a comprehensive overview of the computational details used for bulk and substrate calculations. It includes the optimized bulk lattice parameters, calculated surface energies corresponding to the most stable surfaces and terminations, and the number of substrate layers used. For pristine phosphorene, the calculated lattice parameters and surface energy are consistent with ref. [dhanabalan2017emerging]. The surface energies of Al(111), Cu(111), and Fe(110) agree with previous DFT calculations [jain2013commentary, kiejna2001first, vitos1998surface]. For $\mathrm{Al_{2}O_{3}}$(0001), both Al-terminated and O-terminated configurations were benchmarked against refs. [chiang2016dft, ramogayana2021density, sun2006structure]. The $\mathrm{Cu_{2}O}$(111) O-terminated surface structure and energetics are consistent with refs. [zhang2018atomistic, kokalj2015density, gao2020surface], while $\mathrm{Fe_{2}O_{3}}$(0001) Fe-terminated and O-terminated surfaces were compared with refs. [huang2016surface, nguyen2013water, rohrbach2004ab]. The Cr(110) surface energy matches literature values [jain2013commentary, vitos1998surface], as does the Cr-terminated $\mathrm{Cr_{2}O_{3}}$(0001) surface [sun2006structure, rohrbach2004ab, costa2009ab].
@@
For this study, calculations focused solely on monolayer phosphorene, as the primary interest was in determining the capability of the solid lubricant to adhere on different metal substrates. Other computational studies, where bilayers were considered, have addressed the effects on the cleavage and shear strength of the metallic interfaces [losi2023modeling] and interlayer sliding properies [benini2023interaction].
@@
For each metal oxide substrate, two possible surface terminations were considered by cleaving the structures along the z-axis and exposing outermost atomic layers of metal atoms, ”M-term”, or oxygen atoms, ”O-term”. The relaxed geometries of these different surface terminations are shown as SM.
@@
where $E_{\mathrm{sub}}$ and $E_{\mathrm{bulk}}$ are the total energies of the substrate and of the bulk, respectively, n is the ratio between the number of atoms in the substrate and in the bulk, and A is the area of the supercell. For phosphorene, the monolayer was considered for the calculation of $E_{\mathrm{sub}}$. For oxidized phosphorene, the bulk structure was constructed by replicating the most stable oxidized monolayer configuration along the out-of-plane direction. $E_{\mathrm{bulk}}$ was obtained from this multilayer oxidized structure, ensuring consistency with the oxidation pattern of the monolayer system.
@@
To quantify the structural differences between the interacting structures in the supercell and their isolated counterparts, the Root Mean Square Deviation (RMSD) was employed through a custom Python script. RMSD, defined as $\sqrt{\frac{\sum_{i}d_{i}^{2}}{n}}$ , where $d_{i}$ represents the distance between each of the n pairs of equivalent atoms in two optimally superimposed structures, is commonly used in protein structure analysis and MD simulations for expressing structural similarity [debefve2021systematic, sharma2014accuracy].
@@
An RMSD value of 0 Å indicates identical structures, with increasing values reflecting greater structural differences [carugo2001normalized]. RMSD values are considered reliable indicators of variability when applied to highly similar structures, providing a quantitative measure of geometrical resemblance [carugo2001normalized]. In this study, visual inspection of the phosphorene-substrate interactions revealed noticeable but not excessive distortions. Therefore, RMSD evaluation was considered an appropriate approximation for quantifying the degree of structural distortion.
To analyze the electronic charge redistribution resulting from the interaction between phosphorene and the substrate, the charge displacement was calculated by subtracting the charge densities of the isolated constituents from that of the interface. The resulting $\Delta\rho(x,y,z)$ was then integrated along the xy-plane to obtain the planar average charge density redistribution, $\Delta\rho(z)$. Finally, the total amount of electronic charge redistributed per surface unit ($Q_{red}$) was computed by integrating the absolute value of $\Delta\rho(z)$ along the z-axis and divided by the supercell area [wolloch2018interfacial].
@@ 3 Results and discussion
For oxide substrates, the adhesion energies are generally less favorable (less negative values) compared to their corresponding pristine metals. This can be attributed to the reduced availability of electronic charge for bonding, as oxides typically have fewer delocalized electrons than metals. Among the oxides, $\mathrm{Fe_{2}O_{3}}$ shows the strongest adhesion, particularly with oxidized phosphorene and oxygen termination. $\mathrm{Cu_{2}O}$ exhibits better adhesion with pristine phosphorene, while $\mathrm{Cr_{2}O_{3}}$ shows similar values for both forms. Notably, while $\mathrm{Al_{2}O_{3}}$ presents comparable adhesion to other oxides with pristine phosphorene, its interaction with oxidized phosphorene is extremely weak.
The relationship between the calculated adhesion energies and the experimental tribological performance reported in ref. [vezzelli2024different] reveals important insights. In the experiments, the best results in terms of coefficient of friction (CoF) reduction were observed for scratched iron, where the superficial oxide layer was most likely removed, while non-scratched iron exhibited poor performance. This suggests that phosphorene-substrate adhesion plays a crucial role in enhancing the lubricating properties of BP, being significantly higher on pristine iron than on oxidized iron (Figure 3). This observation is consistent with findings for other layered materials used as solid lubricants, such as MXenes [marquis2025mxene], and can be explained by considering that higher adhesion of the layers to the substrate prevents their peeling off during the rubbing process. Indeed, BP was found to reduce the CoF of non-scratched aluminum substrate only at the beginning of the tribological test, while after rubbing it reached the value measured for the unlubricated sample [vezzelli2024different], in agreement with the low phosphorene-$\mathrm{Al_{2}O_{3}}$ adhesion calculated in this work (Figure 3). The low phosphorene-$\mathrm{Cu_{2}O}$ adhesion may suggest an irrelevant effect of BP in reducing the CoF of copper, as observed for aluminum. However, the experimental results indicate an even worse effect of BP on copper, which was interpreted as third-body wear effect due to the poor adhesion.
Figure 4 presents the evaluation of RMSD for phosphorene-substrate interactions. Analysis of the data reveals a weak correlation between RMSD values and adhesion energies. In general, oxidized phosphorene exhibits higher RMSD values compared to its pristine counterpart, aligning with its more favorable adhesion energies. This suggests that oxidized phosphorene undergoes greater structural distortion due to the more favorable interaction. The structural changes are predominantly confined to the phosphorene layer, with minimal deformation occurring in the substrate materials. Notably, $\mathrm{Cu_{2}O}$ stands as an exception to this trend.
@@
Figure 5: (a) Total electron charge redistribution per surface unit ($Q_{red}$) for every phosphorene-substrate system; (b) $\Delta\rho(z)$ and PDOS analysis for phosphorene deposited on pristine metals and (c) the corresponding oxides.
The analysis of the electronic properties of the considered interfaces is reported in Figure 5. Panel a) presents the overall assessment of charge redistribution ($Q_{red}$) when phosphorene is deposited on the substrate. In agreement with previous studies, e.g., [losi2023modeling, wolloch2018interfacial], higher interfacial adhesion (i.e., more negative adhesion energies) is associated with higher $Q_{red}$ values. Moreover, $Q_{red}$ follows the same trend as adhesion, increasing from Al to Cu, Fe, and Cr in the case of pristine metals, while it remains very low for oxides, except for O-terminated $\mathrm{Fe_{2}O_{3}}$, where the adhesion is higher than for the other oxides. The profiles of the charge displacement, $\Delta\rho(z)$, reveal that a greater amount of electronic charge is accumulated at the interface when phosphorene is deposited on pristine metals (Figure 5b) than oxides (Figure 5c), attributed to the stronger chemisorption on the former substrates. From the PDOS analysis, an increase in phosphorene p-states within the band gap is observed in some cases after interaction with the substrate. In these cases, phosphorene exhibits metallic character [pan2016monolayer]. In some instances, metallic behavior emerges only for oxidized phosphorene, e.g., Figure 5b inset (d), whereas in other cases it is observed for both pristine and oxidized phosphorene, e.g., insets (j,l). Conversely, in many other cases the insulating character is preserved. Moreover, no strong correlation is found between the emergence of metallic behavior in phosphorene and the $Q_{red}$ values, as evidenced by the comparison between Figure 5b insets (j,l) and the $Q_{red}$ values reported for Cr in Figure 5a.